OpenStudy (anonymous):

Write given equation in standard form, solve equation using quadratic formula: 5x^2=12x-3

OpenStudy (anonymous):

Do you know what standard form is?

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, we rearrange the given equation to fit that form

OpenStudy (anonymous):

okay I got 5x^2 -12x+3=0. Then I factor. But I am not getting anything that factors out.

OpenStudy (anonymous):

\[5x ^{2}-12x+3=0\] Then use the quadratic equation.

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

Ahhh

OpenStudy (anonymous):

Got ya

OpenStudy (anonymous):

the quadratic formula will work on any quadratic equation in standard form. Even the ones you factor...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem. Good luck

OpenStudy (anonymous):

Let me know what you get and i'll check it for you

OpenStudy (anonymous):

Found it. I got 12\[12\pm \sqrt{84} Over 10\]

OpenStudy (anonymous):

Yea, now just simplify your answer if you need to

OpenStudy (anonymous):

Thanks again!